OpenStudy (anonymous):

Can anyone help me figure out how the heck they got this answer? Its science Coulombs Law: F = Kq1 q 2 d = distance between the charged particles q1 = quantity of charge of 1 particle q2 = quantity of charge of the other particle K = Constant equal to 9.0*10^9 N-m^2 / C^2 The problem and the answer they got for it will be in a comment. PLEASE HELP!!!! (:

OpenStudy (anonymous):

A hydrogen atom has one proton in its nucleus, which is surrounded by just one electron. The average distance between a proton and an electron in a hydrogen atom is 5.3 * 10^-11 m. What is the electrical force between a proton and an electron in a hydrogen atom? They said that this would be the answer: F = (9.0 * 10^9 N-m^2 / C^2) * [(1.6 * 10^-19 C) (1.6 * 10^-19 C)] / (5.3 * 10^-11 m ^2) F = 8.2 * 10^-8 N

OpenStudy (anonymous):

OH DAMN

OpenStudy (anonymous):

IKR!! >.<

OpenStudy (anonymous):

heeeeeeeeeeelp? ANybody?

OpenStudy (anonymous):

What chu tryin to find.

OpenStudy (anonymous):

it looks really confusing.

OpenStudy (anonymous):

I'm trying to find the electrical fouce between a proton and an electron in a hydrogen adom >.< So many people viewing... yet so little helping :(

OpenStudy (anonymous):

atom*

OpenStudy (anonymous):

Like..the distance?

OpenStudy (anonymous):

force*

OpenStudy (anonymous):

No like the force

OpenStudy (anonymous):

HOlD ON.

OpenStudy (anonymous):

is it a worksheet?

OpenStudy (anonymous):

q1 * q2 F = k-------- r^2 this is the correct formula

OpenStudy (anonymous):

Easy F=kqq/r^2 So \[F=proton charge*electroncharge/(distance^2)\]

OpenStudy (anonymous):

oops *k

OpenStudy (anonymous):

IK! But did you take a look at the work they said is CORRECT? The answer they gave me? the piece where they showed there work?

OpenStudy (anonymous):

Ivy is it a worksheet????

OpenStudy (anonymous):

Ivy is it a worksheet????

OpenStudy (anonymous):

k=9*10^-9 proton charge=1.6*10^-19 electron charge=1.6*10^-19 and the distance between them is=5.3*10^-11 The worksheet is correct, and their answer is correct just follow their work

OpenStudy (anonymous):

Your just repeating the information i already knew... I was ASKING for you to show the steps please. And nope its a lesson for online schooling >.<

OpenStudy (anonymous):

The information is plug and chug. The electrical force is F=kqq/r^2, like in the given information. You just have to find F by plugging in all the known information. There's only one step here.

OpenStudy (anonymous):

IK! But like... i tried doing the math like i plugged everything in like they said. Being the first step is to [(1.6 * 10^-19 C) (1.6 * 10^-19 C)] Which I am pretty sure is 2.56 * 100^361 C ?..... Then you have to multiply (9.0 * 10^9 N-m^2 / C^2) * giving what i think is 23.04 * 1000^3249 Then you divide? IDK :P

OpenStudy (anonymous):

ivy, i do online schooling aswell.. whats the programs name of your online schooling.?

OpenStudy (anonymous):

@creepy after calculation found yr ans correct ....

OpenStudy (anonymous):

wait nonono when you multiply, you don't multiply the exponents, you add them. So \[(1.6 * 10^-19 C) (1.6 * 10^-19 C)=2.56*10^-38\]

OpenStudy (anonymous):

C^2

OpenStudy (anonymous):

@creepy YR ANS IS CORRECT ...

OpenStudy (anonymous):

@ Hark After calculation i found it to NOT be the answer they gave me ^ i wrote it above This is a lesson not an assignment ^.^

OpenStudy (anonymous):

Wait so PJW, you dont multiply the 10s together? ;o

OpenStudy (anonymous):

nope when you multiply 10 to the power of somethings together, you add the exponents. So 10^-19*10^-19=10^-38

OpenStudy (anonymous):

DO YOU WANT ME TO GIVE YOU THE SOLUTION OR WHAT ??????

OpenStudy (anonymous):

Ah, i seeeeeeee So then you get (2.56 * 10^-38) * (9.0 * 10^9 N-m^2 / C^2) ? Which would be.. 23.04 * 10^-27 N-m^2 / C^2 ?

OpenStudy (anonymous):

What do you mean Hark? I already have the solution i just need to know how to get it SPECIFICALLY. Like PJW is showing me

OpenStudy (anonymous):

Well I am off to eat and cool off. Hope you answered my question when i get back >:( Bbl

OpenStudy (anonymous):

okay here it is q1 x q2 = \[(1.6 x 10^{-19})(1.6x10^{-19}) = 2.56 x 10^{-38}\]

OpenStudy (anonymous):

k*q1*q2 = \[(9.0 x 10^{9})(2.56 x 10^{-38}) = 23.04 x 10^{-29}\]

OpenStudy (anonymous):

d (or r) = \[5.3 x 10^{-11}\]

OpenStudy (anonymous):

alright, keep going.. (:

OpenStudy (anonymous):

so \[d ^{2} = (5.3 x 10^{-11})(5.3 x 10^{-11}) = 28.09 x 10^{-22}\]

OpenStudy (anonymous):

so finally k q1q2 F = --------- d^2 \[23.04 x 10^{-29}\]---------------------- \[28.09 x 10^{-22}\]

OpenStudy (anonymous):

Wow! Thats a lot more work than i thought.. keep goin It says i have to get F = 8.2 * 10^-8 N

OpenStudy (anonymous):

this gives \[0.82 x 10^{-29 +22}\] which gives \[0.82 x 10^{-7}\] Now you hv to put it in scientific notation that is move the decimal to the right by one place To achieve this we multiply it by 1 in the form 10/10 the 10 in the numerator multiplies 0.82 to give 8.2 and the 10 in the denominator goes up as 10^-1 so we finally get \[8.2 x 10^{-8}\]

OpenStudy (anonymous):

I hope this is clear now. You have to use the laws of exponents here \[a ^{m} x a ^{n} = a ^{m+n}\] and \[a ^{m}/a ^{n} = a ^{m-n}\]

OpenStudy (anonymous):

Hey creepy... Is it clear now?????

OpenStudy (anonymous):

Oh! Yea its pretty clear now (: Thanks so fricking much ;D